Aricuzyas:

azmine placed $1500 in a certificate of deposit. She earns $15 each month for the next 5 years. Find the annual simple interest rate for the certificate of deposit. *

darkknight:

P = I(1+r)^t I = initial investment, r = rate t = time in years

darkknight:

yep you are given in months so replace ^t to ^(t/12)

jhonyy9:

- so like a first step you need to know how many month has a year - after this multiplie the number of months in a year by 5 and so you will get the total number of months in these 5 years - this number of months of 5 years multiplie by $15 so in this way will get the total interest rate of 5 years but you need to know just the annual simple interest rate - so for this you just need divide the total interest rate of 5 years by 5 and so in this way will get the right correct answer hope helped understandably easy

jhonyy9:

@darkknight your opinion pls. ? ty

darkknight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) darkknight P = I(1+r)^t I = initial investment, r = rate t = time in years \(\color{#0cbb34}{\text{End of Quote}}\) yea right here that is the simple interest formula so jhonyy "so for this you just need divide the total interest rate of 5 years by 5 and so in this way will get the right correct answer " nah i don't think you divide the interest by the number of years Interest is exponential and increases over time, so no we wouldn't be dividing, it goes against the formula and not how interest works. plz try again @jhonyy9

jhonyy9:

so here you dont wrote inside this formula the ,,n" - not is necessary always ?

jhonyy9:

@darkknight

darkknight:

"Find the annual simple interest rate" this problem is SIMPLE interest rate, since not compounded there is no n in the function, the other problem was different @jhonyy9

jhonyy9:

ty

darkknight:

np+